# Computing in a Faulty Congested Clique

### v1

@@ 4 Layered Circuits as $c$-$\mathsf{Faulty~{}Clique~{}}$Algorithms
An $[N,K,d]_{q}$ code is a mapping $enc:(\mathbb{F}_{q})^{K}\rightarrow(\mathbb{F}_{q})^{N}$, such that the Hamming distance between any two codewords $s_{1},s_{2}\in Image(enc)\subseteq(\mathbb{F}_{q})^{N}$ is at most $d$. By its definition, an $[N,K,d]_{q}$ code can correct up to $d-1$ erasures, that is, given a string $s^{\prime}\in(\mathbb{F}_{q})^{N}$, there is at most one codeword $s\in Image(enc)\subseteq(\mathbb{F}_{q})^{N}$ which equals $s^{\prime}$ up to at most $d-1$ erasures of symbols in $(\mathbb{F}_{q})^{N}$.
@@
Let $n$ and $c$ be some integers. Let $p$ be a prime number and let $k$ be an integer such that the prime power $q=p^{k}$ satisfies $\log q=\log(p^{k})=\Theta(cb\log(n))=\Theta(c\log|\Sigma|)$, then there exists an $[N,K,d]_{q}:=[n,n/c,\frac{c-1}{c}n+1]_{q}$ code.
@@
The existence of a code with these parameters ($d=N-K+1$) is a classical result in coding theory; in particular, since $q\geq n$ (i.e., since the field size is bigger than the length of the code) we have that the classical construction of Reed-Solomon codes suffices. See Section 6.8 of [Lin98] or Section 5.2 of [HP03]. ∎
@@
      (Repeat/Fill out Missing Bingo Cards) If all of the parts in a layer have been checkpointed, then the nodes move on to the next epoch; otherwise, they divide the work of the failed nodes and simulate the computations of the gates in their parts until all are checkpointed. In particular, if $F$ nodes failed so far, and of those $F$, there are $F^{c}$ nodes whose parts are yet to have been checkpointed, the $n-F$ non-faulty nodes evenly divide themselves and simulate those $F^{c}$ parts by rerunning steps (i)-(iii). This is done as follows.
@@
        If $F^{c}>n-F$, then each of the $n-F$ non-faulty nodes is assigned to one failed node whose part has not been checkpointed yet, and repeats steps (i)-(iii) for that part.
@@
        Otherwise, there are $F^{c}\leq n-F$ such parts $P_{\tau,v_{1}},P_{\tau,v_{2}},...,P_{v_{F^{c}}}$. We batch these into sets of size at most $6c$, and let each batch be simulated by multiple nodes, as follows. We denote $F^{c}_{batch}=\max\{\lfloor{F^{c}/3c}\rfloor,1\}$, and $mult=\lfloor{\frac{n-F}{F^{c}_{batch}}}\rfloor$. We denote by $x$ the remainder when dividing $n-F$ by $F^{c}_{batch}$ (thus, $mult=\frac{n-F-x}{F^{c}_{batch}}$). The non-faulty nodes, $w_{0},\dots,w_{n-F}$, are split into $F^{c}_{batch}$ batches:
@@
        $R_{\{v_{\left(F^{c}_{batch}-1\right)3c},v_{\left(F^{c}_{batch}-1\right)3c+1},\dots,v_{F^{c}-1}\}}:=\{w_{(F^{c}-1)mult},\dots,w_{{n-F-x}}\}$
        Notice that since $F^{c}_{batch}$ is defined using the floor of the respective ratio, the size of the last batch may be larger than $3c$, but it is at most $6c$. Further, in case $F^{c}$ is less than $3c$ then there is only one batch, which may also be small, but it is assigned to all non-faulty nodes. For each batch, every node in that batch is assigned to each $v_{j}$ in the batch, and repeats steps (i)-(iii) for the corresponding part $P_{\tau,v_{j}}$.
@@
We prove that $R$-decodability holds by induction, with $R=O(cn^{\zeta})=O(n^{\zeta}$ for all epochs except the last, and $R=O(cn^{\mu})=O(n^{\mu})$ for the last. The base case is straightforward because the computation takes place through quiet rounds: at the end of step (1), each node holds the inputs to its gates. In step (2), each node encodes its input using the code in Lemma 5, and thus $R$-decodability holds with $R=O(c)=O(1)$. For the the inductive hypothesis, assume that epoch $\ell$ has been correctly checkpointed as at most $O(n^{\zeta})$ codewords of an $[n,\frac{n}{c},\frac{c-1}{c}n+1]_{q}$-code for each node. Then, by the inductive hypothesis, we have that a node can, in $O(n^{\zeta})$ rounds, obtain all of the checkpointed data it needs and then perform its local computations, and simulate any failed nodes. Thus, at the end of step 3(iv), which we show below that it indeed checkpoints the computation of all nodes (also faulty nodes), we have that epoch $\ell+1$ is correctly checkpointed as at most $O(n^{\zeta})$ codewords of an $[n,\frac{n}{c},\frac{c-1}{c}n+1]_{q}$-code for each node. Thus, $R$-decodability holds with $R=O(cn^{\zeta})=O(n^{\zeta})$. For the last epoch, we only need to encode the output and therefore we get $R$-decodability with $R=O(cn^{\mu})=O(n^{\mu})$.
@@
It remains to bound the number of attempts needed for step 3(iv), which will also prove that indeed it checkpoints the computation of all nodes (also faulty nodes). For an attempt, recall that we denote by $F$ the number of nodes that failed so far (and thus the number of non-faulty nodes is $n-F$), and by $F^{c}$ the number of parts that are yet to be checkpointed. Let $F^{\prime}$ be the additional number of nodes that fail throughout this attempt. We consider the two possible cases, depending on how $F^{c}$ relates to $n-F$.
@@
In the first case, recall that if $F^{c}>n-F$ then each of the non-faulty nodes is assigned to one failed node whose part has not been checkpointed and repeats its computation. Since there are at least $n/c$ non-faulty nodes even despite the additional $F^{\prime}$ newly failed ones, we have that at least $n/c$ additional tasks get checkpointed in this attempt. Thus, such an attempt can happen at most $c$ times before all $n$ parts are checkpointed for this epoch.
@@
In the second case, we have that $F^{c}\leq n-F$. Recall that each part that is not yet checkpointed is now attempted to be checkpointed by a multiplicity of $mult$ or $mult-1$ nodes, where $mult=\left\lceil\frac{n-F}{F^{c}}\right\rceil\geq 1$. There are two possible sub-cases based on how the value of $F^{\prime}$ relates to the remaining number of allowed failures $F_{remain}=(\frac{c-1}{c})\cdot n-F$. It holds that either $F^{\prime}\geq F_{remain}\cdot\frac{1}{2}$ or $F^{\prime}<F_{remain}\cdot\frac{1}{2}$. The former case can happen at most $\log{((\frac{c-1}{c})\cdot n)}=O(\log{n})$ times before $F_{remain}$ drops to 0.
@@
Thus, suppose that the latter case happens in an attempt. If there is a single batch because $F^{c}<3c$, then all of the batch is performed by all of these nodes, and hence must succeed because at least one of them (at least $n/c$ of them) is non-faulty.
@@
Otherwise, since $x$ is the remainder of dividing $n-F$ by $F^{c}_{batch}$, we have $x\leq F^{c}_{batch}-1$. Thus, the actual number of nodes that are successful in checkpointing the tasks they are now in charge of is at least $n-F-F^{\prime}-x$. Possibly, every $mult=\lfloor{(n-F)/F^{c}_{batch}}\rfloor$ of them are performing the same batch, but even in this worst case, the number of distinct newly checkpointed batches is at least $(n-F-F^{\prime}-x)/mult$. Note that with the notation of $x$, we have that $mult=(n-F-x)/F^{c}_{batch}$. Further, it holds that $F^{c}_{batch}=\max\{\lfloor F^{c}/(3c)\rfloor,1\}\leq\frac{n}{3c}\leq(n-F)/3$, where the first inequality is since $F^{c}<n$, and the second is since $n/c\leq n-F$ because the number of non-faulty nodes can never go below $n/c$.
@@
We bound this number of newly checkpointed batches $(n-F-F^{\prime}-x)/mult$ in terms of the total number of remaining batches $F^{c}_{batch}$, as follows. We have that:
$\frac{n-F-F^{\prime}-x}{mult}$ $=(n-F-F^{\prime}-x)/\frac{n-F-x}{F^{c}_{batch}}$
$=F^{c}_{batch}\cdot\frac{n-F-F^{\prime}-x}{n-F-x}=F^{c}_{batch}\cdot(1-\frac{F^{\prime}}{n-F-x}).$
@@
$\frac{n-F-F^{\prime}-x}{mult}$ $\geq F^{c}_{batch}\cdot(1-\frac{F_{remain}\cdot\frac{1}{2}}{n-F-x})=F^{c}_{batch}\cdot(1-\frac{((\frac{c-1}{c})\cdot n-F)\cdot\frac{1}{2}}{n-F-x}).$
@@
$\frac{n-F-F^{\prime}-x}{mult}$ $\geq F^{c}_{batch}\cdot\left(1-\frac{n-F-n/c}{2(n-F-x)}\right)$
$=F^{c}_{batch}\cdot\left(1-\frac{(n-F)-n/c}{2(n-F)-2x}\right)=F^{c}_{batch}\cdot\left(1-\frac{1-\frac{n/c}{n-F}}{2-\frac{2x}{n-F}}\right).$
Removing the $\frac{n/c}{n-F}$ term from the nominator only decreases the value of the expression, as well as using $x\leq F^{c}_{batch}-1$. We thus have:
$\frac{n-F-F^{\prime}-x}{mult}$ $\geq F^{c}_{batch}\cdot\left(1-\frac{1}{2-\frac{2(F^{c}_{batch}-1)}{n-F}}\right).$
Since $\frac{n-F}{F^{c}_{batch}}\geq 3$, we finally have:
$\frac{n-F-F^{\prime}-x}{mult}$ $\geq F^{c}_{batch}\cdot\left(1-\frac{1}{2-\frac{2(F^{c}_{batch}-1)}{3F^{c}_{batch}}}\right)\geq F^{c}_{batch}\cdot\left(1-\frac{1}{2-\frac{2}{3}}\right)=F^{c}_{batch}\cdot(1/4).$
This yields that after this subcase, the number of batches to be checkpointed drops from $F^{c}_{batch}$ to at most $(1/4)\cdot F^{c}_{batch}$. This implies that this can happen at most $O(\log{n})$ times before all $n$ parts are checkpointed.
@@ 4.2 Coding is Essential
Consider a function $f^{\mathcal{A}}$ computed by a clique algorithm $\mathcal{A}$, and assume that it is incompressible for the string that represents the input of a node $w$. Then, if there are $c$ quiet rounds and they form a mega-round, then node $w$ must locally encode all of their data using an MDS (not necessarily linear) code of type $[n,n/c,(c-1)n/c+1]_{q}$.
@@
We prove that if $w$ does not encode all of its data then, the rest of the network is not able to distinguish between two of $w$’s states, which contradicts the function $f^{\mathcal{A}}$ being incompressible. Consider the messages that node $w$ sends to all of the nodes in the first $c$ quiet rounds as a word $s_{w}\in\Sigma^{cn}$, which can be considered a codeword of an $[n,\frac{n}{c},d]_{q}$-code. The proof is completed by showing that the distance of this code must be $d=\frac{c-1}{c}n+1$. Suppose that the distance $d$ of the code is at most $D=\frac{c-1}{c}n$. Then there is an input $x_{w}$ for which the Hamming ball around its corresponding codeword $B_{D}(s_{w})$ contains the codeword $s_{w}^{\prime}$ of another input $x_{w}^{\prime}\neq x_{w}$, so that $s_{w}$ cannot be uniquely decoded (or distinguished from $s_{w}^{\prime}$) given $D$ failures. If the adversary fails $w$ along with $\frac{c-1}{c}n-1$ other nodes in the first noisy round, then considering $s_{w}$ as a codeword we have more than $d$ erasures in it. This means that there are two states (i.e., inputs) of $w$ that are indistinguishable for the rest of the network, so if the nodes correctly compute $f^{\mathcal{A}}$, this contradicts the assumption that $f^{\mathcal{A}}$ is incompressible. ∎

### v2

@@ 4 Layered Circuits as $c$-$\mathsf{Faulty~Clique~}$Algorithms
An $N,K,d]_{q}$ code is a mapping $enc:(\mathbb{F}_{q})^{K}\rightarrow(\mathbb{F}_{q})^{N}$, such that the Hamming distance between any two codewords $s_{1},s_{2}\in Image(enc)\subseteq(\mathbb{F}_{q})^{N}$ is at least $d$. By its definition, an $N,K,d]_{q}$ code can correct up to d1d-1 erasures, that is, given a string $s^{\prime}\in(\mathbb{F}_{q})^{N}$, there is at most one codeword $s\in Image(enc)\subseteq(\mathbb{F}_{q})^{N}$ which equals $s^{\prime}$ up to at most d1d-1 erasures of symbols in $(\mathbb{F}_{q})^{N}$.
@@
Let $n$ and $c$ be some integers. Let $p$ be a prime number and let $k$ be an integer such that the prime power $q=p^{k}$ satisfies $\log q=\log(p^{k})=\Theta(cb\log(n))=\Theta(c\log|\Sigma|)$, then there exists an $N,K,d]_{q}:=[n,n/c,\frac{c-1}{c}n+1]_{q}$ code.
@@
The existence of a code with these parameters (d=NK+1d=N-K+1) is a classical result in coding theory; in particular, since $q\geq n$ (i.e., since the field size is bigger than the length of the code) we have that the classical construction of Reed-Solomon codes suffices. See Section 6.8 of [Lin98] or Section 5.2 of [HP03]. ∎
@@
      (Repeat/Fill out Missing Bingo Cards) If all of the parts in a layer have been checkpointed, then the nodes move on to the next epoch; otherwise, they divide the work of the failed nodes and simulate the computations of the gates in their parts until all are checkpointed. In particular, if $F$ nodes failed so far, and of those $F$, there are $F^{*}$ nodes whose parts are yet to have been checkpointed, the nFn-F non-faulty nodes evenly divide themselves and simulate those $F^{*}$ parts by rerunning steps (i)-(iii). This is done as follows.
@@
        If $F^{*}>n-F$, then each of the nFn-F non-faulty nodes is assigned to one failed node whose part has not been checkpointed yet, and repeats steps (i)-(iii) for that part.
@@
        Otherwise, there are $F^{*}\leq n-F$ such parts $P_{\tau,v_{1}},P_{\tau,v_{2}},...,P_{v_{F^{*}}}$. We batch these into sets of size at most $6c$, and let each batch be simulated by multiple nodes, as follows. We denote $F^{*}_{batch}=\max\{\lfloor{F^{*}/3c}\rfloor,1\}$, and $mult=\lfloor{\frac{n-F}{F^{*}_{batch}}}\rfloor$. We denote by $x$ the remainder when dividing nFn-F by $F^{*}_{batch}$ (thus, $mult=\frac{n-F-x}{F^{*}_{batch}}$). The non-faulty nodes, $w_{0},\dots,w_{n-F}$, are split into $F^{*}_{batch}$ batches:
@@
        $R_{\{v_{\left(F^{*}_{batch}-1\right)3c},v_{\left(F^{*}_{batch}-1\right)3c+1},\dots,v_{F^{*}-1}\}}:=\{w_{(F^{*}-1)mult},\dots,w_{{n-F-x}}\}$
        Notice that since $F^{*}_{batch}$ is defined using the floor of the respective ratio, the size of the last batch may be larger than $3c$, but it is at most $6c$. Further, in case $F^{*}$ is less than $3c$ then there is only one batch, which may also be small, but it is assigned to all non-faulty nodes. For each batch, every node in that batch is assigned to each $v_{j}$ in the batch, and repeats steps (i)-(iii) for the corresponding part $P_{\tau,v_{j}}$.
@@
We prove that $R$-decodability holds by induction, with $R=O(cn^{\zeta})=O(n^{\zeta}$ for all epochs except the last, and $R=O(cn^{\mu})=O(n^{\mu})$ for the last. The base case is straightforward because the computation takes place through quiet rounds: at the end of step (1), each node holds the inputs to its gates. In step (2), each node encodes its input using the code in Lemma 6, and thus $R$-decodability holds with $R=O(c)=O(1)$. For the the inductive hypothesis, assume that epoch $\ell$ has been correctly checkpointed as at most $O(n^{\zeta})$ codewords of an $n,\frac{n}{c},\frac{c-1}{c}n+1]_{q}-$code for each node. Then, by the inductive hypothesis, we have that a node can, in $O(n^{\zeta})$ rounds, obtain all of the checkpointed data it needs and then perform its local computations, and simulate any failed nodes. Thus, at the end of step 3(iv), which we show below that it indeed checkpoints the computation of all nodes (also faulty nodes), we have that epoch $\ell+1$ is correctly checkpointed as at most $O(n^{\zeta})$ codewords of an $n,\frac{n}{c},\frac{c-1}{c}n+1]_{q}-$code for each node. Thus, $R$-decodability holds with $R=O(cn^{\zeta})=O(n^{\zeta})$. For the last epoch, we only need to encode the output and therefore we get $R$-decodability with $R=O(cn^{\mu})=O(n^{\mu})$.
@@
It remains to bound the number of attempts needed for step 3(iv), which will also prove that indeed it checkpoints the computation of all nodes (also faulty nodes). For an attempt, recall that we denote by $F$ the number of nodes that failed so far (and thus the number of non-faulty nodes is nFn-F), and by $F^{*}$ the number of parts that are yet to be checkpointed. Let $F^{\prime}$ be the additional number of nodes that fail throughout this attempt. We consider the two possible cases, depending on how $F^{*}$ relates to nFn-F.
@@
In the first case, recall that if $F^{*}>n-F$ then each of the non-faulty nodes is assigned to one failed node whose part has not been checkpointed and repeats its computation. Since there are at least $n/c$ non-faulty nodes even despite the additional $F^{\prime}$ newly failed ones, we have that at least $n/c$ additional tasks get checkpointed in this attempt. Thus, such an attempt can happen at most $c$ times before all $n$ parts are checkpointed for this epoch.
@@
In the second case, we have that $F^{*}\leq n-F$. Recall that each part that is not yet checkpointed is now attempted to be checkpointed by a multiplicity of multmult or mult1mult-1 nodes, where $mult=\left\lceil\frac{n-F}{F^{*}}\right\rceil\geq 1$. There are two possible sub-cases based on how the value of $F^{\prime}$ relates to the remaining number of allowed failures $F_{remain}=(\frac{c-1}{c})\cdot n-F$. It holds that either $F^{\prime}\geq F_{remain}\cdot\frac{1}{2}$ or $F^{\prime}<F_{remain}\cdot\frac{1}{2}$. The former case can happen at most $\log{((\frac{c-1}{c})\cdot n)}=O(\log{n})$ times before $F_{remain}$ drops to 0.
@@
Thus, suppose that the latter case happens in an attempt. If there is a single batch because $F^{*}<3c$, then all of the batch is performed by all of these nodes, and hence must succeed because at least one of them (at least $n/c$ of them) is non-faulty.
@@
Otherwise, since $x$ is the remainder of dividing nFn-F by $F^{*}_{batch}$, we have $x\leq F^{*}_{batch}-1$. Thus, the actual number of nodes that are successful in checkpointing the tasks they are now in charge of is at least $n-F-F^{\prime}-x$. Possibly, every $mult=\lfloor{(n-F)/F^{*}_{batch}}\rfloor$ of them are performing the same batch, but even in this worst case, the number of distinct newly checkpointed batches is at least $(n-F-F^{\prime}-x)/mult$. Note that with the notation of $x$, we have that $mult=(n-F-x)/F^{*}_{batch}$. Further, it holds that $F^{*}_{batch}=\max\{\lfloor F^{*}/(3c)\rfloor,1\}\leq\frac{n}{3c}\leq(n-F)/3$, where the first inequality is since $F^{*}<n$, and the second is since $n/c\leq n-F$ because the number of non-faulty nodes can never go below $n/c$.
@@
We bound this number of newly checkpointed batches $(n-F-F^{\prime}-x)/mult$ in terms of the total number of remaining batches $F^{*}_{batch}$, as follows. We have that:
$\frac{n-F-F^{\prime}-x}{mult}$ $=(n-F-F^{\prime}-x)/\frac{n-F-x}{F^{*}_{batch}}$
$=F^{*}_{batch}\cdot\frac{n-F-F^{\prime}-x}{n-F-x}=F^{*}_{batch}\cdot(1-\frac{F^{\prime}}{n-F-x}).$
@@
$\frac{n-F-F^{\prime}-x}{mult}$ $\geq F^{*}_{batch}\cdot(1-\frac{F_{remain}\cdot\frac{1}{2}}{n-F-x})=F^{*}_{batch}\cdot(1-\frac{((\frac{c-1}{c})\cdot n-F)\cdot\frac{1}{2}}{n-F-x}).$
@@
$\frac{n-F-F^{\prime}-x}{mult}$ $\geq F^{*}_{batch}\cdot\left(1-\frac{n-F-n/c}{2(n-F-x)}\right)$
$=F^{*}_{batch}\cdot\left(1-\frac{(n-F)-n/c}{2(n-F)-2x}\right)=F^{*}_{batch}\cdot\left(1-\frac{1-\frac{n/c}{n-F}}{2-\frac{2x}{n-F}}\right).$
Removing the $\frac{n/c}{n-F}$ term from the nominator only decreases the value of the expression, as well as using $x\leq F^{*}_{batch}-1$. We thus have:
$\frac{n-F-F^{\prime}-x}{mult}$ $\geq F^{*}_{batch}\cdot\left(1-\frac{1}{2-\frac{2(F^{*}_{batch}-1)}{n-F}}\right).$
Since $\frac{n-F}{F^{*}_{batch}}\geq 3$, we finally have:
$\frac{n-F-F^{\prime}-x}{mult}$ $\geq F^{*}_{batch}\cdot\left(1-\frac{1}{2-\frac{2(F^{*}_{batch}-1)}{3F^{*}_{batch}}}\right)\geq F^{*}_{batch}\cdot\left(1-\frac{1}{2-\frac{2}{3}}\right)=F^{*}_{batch}\cdot(1/4).$
This yields that after this subcase, the number of batches to be checkpointed drops from $F^{*}_{batch}$ to at most $(1/4)\cdot F^{*}_{batch}$. This implies that this can happen at most $O(\log{n})$ times before all $n$ parts are checkpointed.
@@ 4.2 Coding is Essential
Consider a function $f^{\mathcal{A}}$ computed by a clique algorithm $\mathcal{A}$, and assume that it is incompressible for the string that represents the input of a node $w$. Then, if there are $c$ quiet rounds and they form a mega-round, then node $w$ must locally encode all of their data using an MDS (not necessarily linear) code of type $n,n/c,(c-1)n/c+1]_{q}$.
@@
We prove that if $w$ does not encode all of its data then, the rest of the network is not able to distinguish between two of $w$’s states, which contradicts the function $f^{\mathcal{A}}$ being incompressible. Consider the messages that node $w$ sends to all of the nodes in the first $c$ quiet rounds as a word $s_{w}\in\Sigma^{cn}$, which can be considered a codeword of an $n,\frac{n}{c},d]_{q}-$code. The proof is completed by showing that the distance of this code must be $d=\frac{c-1}{c}n+1$. Suppose that the distance $d$ of the code is at most $D=\frac{c-1}{c}n$. Then there is an input $x_{w}$ for which the Hamming ball around its corresponding codeword $B_{D}(s_{w})$ contains the codeword $s_{w}^{\prime}$ of another input $x_{w}^{\prime}\neq x_{w}$, so that $s_{w}$ cannot be uniquely decoded (or distinguished from $s_{w}^{\prime}$) given $D$ failures. If the adversary fails $w$ along with $\frac{c-1}{c}n-1$ other nodes in the first noisy round, then considering $s_{w}$ as a codeword we have more than $d$ erasures in it. This means that there are two states (i.e., inputs) of $w$ that are indistinguishable for the rest of the network, so if the nodes correctly compute $f^{\mathcal{A}}$, this contradicts the assumption that $f^{\mathcal{A}}$ is incompressible. ∎